### v3

# Description of DG–algebras structure via characters on a groupoid

Abstract
We construct a description of graded derivations in group algebras. Using this result for arbitrary grading of the group algebra, we describe all possible DG–algebras structures. Examples are given. The description is given in terms of characters on a groupoid analogous to the groupoid of conjugate action.
Introduction
In several previous papers bye A.S. Mishchenko and others [AMS17, AM19, AAS23], was presented a method for studying the derivations of group algebras using the character space on the groupoid of conjugacy action. In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebras and all possible DG–algebra structures on group algebras. [^1] The results of Sections 1.1 and 1.2 were obtained by A. Arutyunov, the other results were obtained jointly. The work of A. Arutyunov was supported by the RSF grant No. 25-11-00018.
DG–algebras and graded derivations are occur frequently. See e.g. [Orl23], where they are studied in the context of algebraic geometry. Graded derivations in isolation are an even better known subject, see e.g. [Mei03]. The classification of possible DG structures has also been a focus of study [Mao11, MWZ22].
The main object of our study will be the graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivations, that is linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ satisfying the graded Leibniz rule
$d(uv)=d(u)v+(-1)^{|u|}ud(v),\quad\forall u,v\in A_{i}.$
In Section 1 we present the structure of a groupoid which is connected to the graded algebra structure. We will give a description of graded derivations as locally finite characters on this groupoid (see Theorem 1.1). In Section 2, we describe the inner, quasi-inner and central derivations in the graded case. Using this construction in the case of DG–algebras, in Section 3 we obtain a description of DG–algebra structures in terms of characters and an isomorphism criterion for DG–algebras. In addition, at the end of Section 3, we give several examples of gradings on group algebras and derivations that define a DG–algebra structure on them. The definitions and properties from combinatorial group theory that are necessary to understand the present paper can be found in [LS01]
1 The groupoid, it’s characters, and graded derivations.
Let’s start with the definitions and fix the notation for future reference. We will study derivations of group algebras with a DG structure. The standard definition of a group algebra is
Definition 1.1.
Let $G$ be a group and let $K$ a field. The group algebra $K[G]$ is the associative algebra over $K$ the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are defined by:
$\sum_{g\in G}a_{g}g+\sum_{g\in G}b_{g}g=\sum_{g\in G}(a_{g}+b_{g})g,$
$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}\Big{(}\sum_{xy=h\in G}(a_{x}b_{y})h\Big{)}.$
The elements of the group $G$ form a basis of $K[G]$. Multiplication of the basis elements in the group algebra is induced by the group multiplication.
Furthermore, we need the notion of a groupoid. usually, it is defined as a category in which every morphism is invertible. We will work with the following construction associated with a finitely generated group $G$, so this category is small (the classes of objects and morphisms are sets).
Definition 1.2.
A groupoid $\Gamma$ has a set of objects $\operatorname{Obj}(\Gamma)$ and a set of morphisms $\operatorname{Hom}(\Gamma)$. Each morphism has a source $s(\phi)$ and a target $t(\phi)$, i.e. each morphism is in the set $\operatorname{Hom}(s(\phi),t(\phi))$. Composable morphisms have an associative operation (composition). The following conditions hold:
  1. 1.
    For each object $a\in\operatorname{Obj}(\Gamma)$, there is a neutral endomorphism $1_{a}\in\operatorname{Hom}(a,a)$ such that $\forall\phi\in\operatorname{Hom}(b,a)$: $1_{a}\circ\phi=\phi$ and $\forall\phi\in\operatorname{Hom}(a,b)$ $\phi\circ 1_{a}=\phi$.
  2. 2.
    Each morphism has the inverse: fir every $\phi\in\operatorname{Hom}(a,b)$ there exists $\psi\in\operatorname{Hom}(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
1.1 Groupoid of conjugacy action.
The following construction generalizes the one from [AMS17]. Suppose we have a $\mathbb{C}$–graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$. The group $G$ is assume to be finitely generated. We will define the groupoid of conjugacy action $\Gamma$ for the group $G$ (in the sense of Definition 1.2) as follows:
[…]
Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition $\psi\circ\phi$ should be defined. Define it as follows:
$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1}).$
This formula is conveniently depicted as a diagram.
$(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=u_{2}v_{2}^{-1}$$u_{1}v_{1}^{-1}$$(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$$\scriptstyle{(u_{2},v_{2})}$$\scriptstyle{(u_{1},v_{1})}$$\scriptstyle{(u_{2},v_{2})\circ(u_{1},v_{1})}$
Proposition 1.1.
$\Gamma$ is a groupoid.
Proof.
First of all, we note that $\exists g\rightarrow h\Leftrightarrow\exists t:h=(-1)^{|t|}tgt^{-1}.$
Consider $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Define $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
We need to check that
$s(\psi\circ\phi)=a$ and $t(\psi\circ\phi)=c$.
By definition,
$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|\mathstrut}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}$ = $(-1)^{|v_{1}|\mathstrut}v_{1}^{-1\mathstrut}u_{1}=a$.
Calculate that
$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}=u_{2}v_{2}^{-1\mathstrut}=c$, because $u_{1}v_{1}^{-1\mathstrut}=(-1)^{|v_{2}|\mathstrut}v_{2}^{-1\mathstrut}u_{2}$
Here we shall prove that the composition is associative. Let $\phi_{1},\phi_{2},\phi_{3}$ be morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$. We need to show that
$\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(\phi_{1}\circ\phi_{2})\circ\phi_{3}.$
We have proved above that these compositions are well–defined. The left side is
$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
While the right one is equal to it,
$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
proving associativity.
Let us show that there is an element $1_{g}$ such that
$s(1_{g})=t(1_{g})=g$,
$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $\operatorname{Hom}(\Gamma).$
Take $1_{g}=(g,e)$, where $e$ is the identity of the group $G$.
$(-1)^{|e|}e^{-1}g=g=ge^{-1}$$\scriptstyle{(g,e)}$
Indeed, choosing an arbitrary morphism $(u,v)$, we have:
$(uv^{-1},e)(u,v)=(u,v)$;
$(u,v)((-1)^{|v|}v^{-1}u,e)=((-1)^{|v|}v(-1)^{|v|}v^{-1}u,e)=(u,v)$.
We have now shown that $\Gamma$ is a category. It remains to check the invertibility of the arrows.
We claim the the inverse of a morphism of the form $(u,v)$ is $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
Indeed,
$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e);=(t(\phi),e).$
$((-1)^{|v|}v^{-1}uv^{-1},v^{-1})(u,v)=((-1)^{|v|}v^{-1}u,e)=(s(\phi),e).$
$(-1)^{|v|}v^{-1}u$$uv^{-1}$$\scriptstyle{(u,v)}$$\scriptstyle{((-1)^{|v|}v^{-1}uv^{-1},v^{-1})}$
This proves Proposition 1.1. ∎
The following properties of the groupoid are easily verified by direct calculation.
Proposition 1.2.
  1. 1.
    The endomorphisms $\operatorname{Hom}(a,a),a\in\pm G$, admit the following description:
    $\operatorname{Hom}(a,a)=\{\chi\in\operatorname{Hom}(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
  2. 2.
    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$, denote by $\Gamma_{[a]}$ the subgroupoid such that $\operatorname{Obj}(\Gamma_{[a]})=[a]$. The following properties hold:
    1. (a)
      $\operatorname{Hom}(\Gamma_{[a]})=\{\phi\in\operatorname{Hom}(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
    2. (b)
      $\Gamma=\coprod\Gamma_{[a]}$.
  3. 3.
    The left action $\operatorname{Hom}(a,a)\times\operatorname{Hom}(a,b)\rightarrow\operatorname{Hom}(a,b)$: $(\phi,\psi)\to\psi\circ\phi$ is free and transitive.
Let us show that, with the help of characters of $\Gamma$, it is possible to give a description of derivations satisfying the graded Leibniz rule.
Definition 1.3.
The linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
$d(uv)=d(u)v+(-1)^{|u|}ud(v)$, $\forall u,v\in A_{i}$
Definition 1.4.
The map $\chi$ : $\operatorname{Hom}(\Gamma)$ $\to$ $\mathbb{C}$ is a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$:
$\bullet$$\bullet$$\bullet$$\scriptstyle{\chi(\psi)}$$\scriptstyle{\chi(\phi)}$$\scriptstyle{\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)}$
Definition 1.5.
A character $\chi$ is locally finite if for all $v\in\pm G$ we have $\chi(u,v)=0$ for all except a finite number of $u$.
We will denote by $\chi(\Gamma)$ the space of locally finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural Lie algebra structure.
The graded Leibniz rule in a group algebra can be expressed in terms of characters on the groupoid $\Gamma$ as follows.
Proposition 1.3.
For any graded derivation $d$ there exists a locally finite character $\chi$ such that
$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g;\forall u\in A_{i}.$ (1)
Proof.
It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$ as a vector space.
For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$.
Let us define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and show that $\chi$ it is a character. We have
$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}),$
or after a basis expansion
$\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{h}hg_{1}+(-1)^{|g_{2}|}g_{2}\sum\limits_{h\in G}d_{g_{1}}^{h}h.$
It is clear that summing over all $h$ and all $gh$ is equivalent, so we can convert the right-hand side to the following form:
$\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{hg_{1}^{-1}}h+\sum\limits_{h\in G}d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}h\Rightarrow d_{g_{2}g_{1}}^{h}=d_{g_{2}}^{hg_{1}^{-1}}+d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}.$ (2)
Let’s make a substitution:
$h_{1}=hg_{1}^{-1},$
$h_{2}=(-1)^{|g_{2}|}g_{2}^{-1}h.$
Then the last equality (2) takes the form:
$d_{g_{2}g_{1}}^{(-1)^{|g_{2}|}g_{2}h_{2}}=d_{g_{2}}^{h_{1}}+d_{g_{1}}^{h_{2}}$ $\Leftrightarrow$ $\chi(\phi((-1)^{|g_{2}|}g_{2}h_{2},g_{2}g_{1}))=\chi(\phi(h_{1},g_{2}))+\chi(\phi(h_{2},g_{1})).$
It remains only to note that $((-1)^{|g_{2}|}g_{2}h_{2},g_{2}g_{1})=(h_{1},g_{2})(h_{2},g_{1}).$
A corollary of the above statement is the following theorem, proved similarly to the non-graded case, see Theorem 1 of [Aru20].
Theorem 1.1.
For any graded derivation $d$ there exists a unique locally finite character $\chi_{d}$ satisfying
$d(a)=(-1)^{|a|}a\Big{(}\sum\limits_{t\in G}\chi((-1)^{|a|}at,a)t\Big{)}.$
There is therefore an isomorphism $\operatorname{Der}(\mathbb{C}[G])\cong\chi(\Gamma)$.
Proof.
To prove this, we need to apply formula (1) to $a\in\pm G$, write an arbitrary element $g\in G$ as $g=at$ and then drop the terms with a zero coefficient from the sum. See the proof of Theorem 1 in [Aru20] for more details. ∎
1.2 Characters.
Locally finite characters are isomorphic to derivations, which means that the commutator of derivations yields a binary operation on characters. Formula (1) allows us to define a commutator in character space. Let $d_{1},d_{2}$ be the derivations defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$. Denote the character that corresponding to the commutator $d_{1},d_{2}]$ by $\chi_{[d_{1},d_{2}]}$. This defines an operation on the space of characters.
$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}.$ (3)
For this operation, we have the following equality:
Proposition 1.1.
The operation on characters above can be expressed as
$\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g).$ (4)
Proof.
Let $g\in G$. Then we have
$d_{1}(g)$ $=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)h,$
$d_{2}(g)$ $=\sum\limits_{h\in G}\chi_{d_{2}}(h,g)h,$
$\{\chi_{d_{1}},\chi_{d_{2}}\}(g)$ $=\sum\limits_{h\in G}\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)a.$
Write down the equation for the commutator
$d_{1},d_{2}]=d_{1}d_{2}-d_{2}d_{1}$,
$d_{1}d_{2}(g)=\sum\limits_{h\in G}\chi_{d_{2}}(h,g)\Big{(}\sum\limits_{a\in\pm G}\chi_{d_{1}}(a,h)a\Big{)}$,
$d_{2}d_{1}(g)=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)\Big{(}\sum\limits_{a\in\pm G}\chi_{d_{2}}(a,h)a\Big{)}$.
Thanks to local finiteness, we can swap the sums in the last equation and get
$d_{1},d_{2}](h)=\sum\limits_{a\in G}\Big{(}\sum\limits_{h\in G}\chi_{d_{2}}(h,g)\chi_{d_{1}}(a,h)-\chi_{d_{1}}(h,g)\chi_{d_{2}}(a,h)\Big{)}a$.
The value of $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)$ is the coefficient of $a$, so
$\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g)$.
It follows that the $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra, isomorphic to the algebra of all graded derivations by proposition 1.3.
2 Examples of graded derivations and their description in terms of characters.
We will illustrate the correspondence between characters on a groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations. They will serve as easy examples of such operators.
2.1 Inner derivations
Consider an element $a\in\pm G$. Define the mapping $\chi^{a}:\operatorname{Hom}(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming $a\neq b$):
$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in\operatorname{Hom}(b,a)$}\\ -1,&\text{$\phi\in\operatorname{Hom}(a,b)$}\\ 0,&\text{otherwise.}\end{cases}$ (5)
Clearly $\chi^{a}$ is trivial on endomorphisms.
Proposition 2.1.
The map $\chi^{a}$ defines an inner derivations for $a\in\pm G$ by the formula
$d_{a}:x\mapsto[a,x]_{grad}=ax-(-1)^{|x|}xa$.
Proof.
Let’s check that $d_{a}$ is a graded derivation (that is, satisfies the graded Leibniz rule):
$d_{a}(x)y+(-1)^{|x|}xd_{a}(y)=(ax-(-1)^{|x|}xa)y+(-1)^{|x|}x(ay-(-1)^{|y|}ya)=axy-(-1)^{|xy|}xya=d_{a}(xy)$.
Let’s check that it really sets the necessary derivation. Due to linearity, it is enough to check the statement on the element $g\in\pm G$. According to the previous corollary, we have an equality:
$d_{a}(g)=\sum\limits_{h\in G}\chi(h,g)h.$
We have that $\chi(h,g)\neq 0\Leftrightarrow t((h,g))=a$ or $s((h,g))=a$. Let’s consider these two cases:
  1. 1.
    $s((h,g))=a\Leftrightarrow(-1)^{|g|}g^{-1}h=a\Leftrightarrow h=(-1)^{|g|}ga.$
  2. 2.
    $t((h,g))=a\Leftrightarrow hg^{-1}=a\Leftrightarrow h=ag.$
Therefore, we have: $d_{a}(g)=ag-(-1)^{|g|}ga.$
2.2 Central derivations
We will now give an example of a derivation that is not inner. We will show that it admits an simple explicit form and is not trivial on endomorphisms. That entails that it cannot be expressed as a (perhaps formal) sum of inner derivations.
Definition 2.1.
That the map $\tau:G\rightarrow\mathbb{C}$ is a graded group character if it satisfies the following property:
$\tau(ab)=\tau(a)+(-1)^{|a|}\tau(b),$ $\forall a,b\in G$.
Let’s fix a graded group character $\tau$, and let $z\in Z(G)$ be central element of the group.
Definition 2.2.
Define the map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
$d_{z}^{\tau}:g\rightarrow\tau(g)gz$
and extend it by linearity to the whole algebra $\mathbb{C}[G]$. We will call it a central graded derivation.
Proposition 2.2.
The map $d_{z}^{\tau}$ is a graded derivation.
Proof.
It is enough to check the graded Leibniz rule on arbitrary generators $u,v\in G$:
$d_{z}^{\tau}(uv)=\tau(uv)uvz=\tau(u)uzv+(-1)^{|u|}u\tau(v)vz=d_{z}^{\tau}(u)v+(-1)^{|u|}ud_{z}^{\tau}(v)$.
On arbitrary elements of a group algebra, the graded Leibniz rule is satisfied due to the linearity of $d_{z}^{\tau}$. ∎
Proposition 2.3.
Nontrivial central derivations are not inner.
Proof.
Let $\chi_{\tau,z}$ be a character that corresponds to the derivation $d_{z}^{\tau}$. By corollary 1 we have the equality
$d_{z}^{\tau}(g)=\tau(g)gz=g\Big{(}\sum\limits_{t\in G}\chi((-1)^{|g|}gt,g)(-1)^{|g|}t\Big{)}$.
Therefore, this character is not equal to 0 only in the case $t=z$:
$\chi((-1)^{|g|}gz,g)=(-1)^{|g|}\tau(g)$.
Define the morphism $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
$s(\phi)=z$,
$t(\phi)=(-1)^{|g|}z.$
Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ is non–trivial on a loop $\phi$, thus $d_{z}^{\tau}$ it is not an inner derivation. ∎
A direct check shows that inner and central derivations constitute subalgebras. For more details see [Aru20].
2.3 Quasi-inner derivations
As seen from the formula (5), inner derivations are trivial on endomorphisms; however, the converse is not true. We will call such derivations quasi-inner.
Definition 2.3.
Define $\operatorname{QInnDer}(\mathbb{C}[G])$ as follows:
$\operatorname{QInnDer}(\mathbb{C}[G])=\left\{d\in\operatorname{Der}\mid\forall a\in\operatorname{Obj}(\Gamma),\forall\phi\in\operatorname{Hom}(a,a):\chi_{d}(\phi)=0\right\}.$ (6)
Quasi-inner derivations are clearly a subspace, but the more significant statement is that they constitute an ideal in the algebra of all derivations.
Proposition 2.4.
$\operatorname{QInnDer}(\mathbb{C}[G])$ $\triangleleft$ $\operatorname{Der}(\mathbb{C}[G])$ is an ideal.
Proof.
Let us prove that $\operatorname{QInnDer}(\mathbb{C}[G])$ is a subalgebra, i.e.
$d_{1}$, $d_{2}$ $\in$ $\operatorname{QInnDer}(\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in\operatorname{QInnDer}(\mathbb{C}[G])$.
The character $\chi_{d_{1}}$ can be represented as
$\chi_{d_{1}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$, $\lambda^{a}\in\mathbb{C}$,
where $\chi_{d_{1}}$ defined by formula (5), since $\chi^{a}$ is trivial on endomorphisms. Similarly for the character $\chi_{d_{2}}$:
$\chi_{d_{2}}=\sum\limits_{b\in G}\lambda^{b}\chi^{b}$, $\lambda^{b}\in\mathbb{C}$.
Because of the bilinearity of the commutator,
$\{\chi_{d_{1}},\chi_{d_{2}}\}=\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\{\chi^{a},\chi^{b}\}$.
The commutator of the $d^{a}$ и $d^{b}$ can be represented as follows:
$d^{a},d^{b}]=d^{ab}-d^{ba}$.
Then define the character $\{\chi^{a},\chi^{b}\}$ by the following formula
$\{\chi^{a},\chi^{b}\}=\chi^{ab}-\chi^{ba}$.
Claim:
$\{\chi_{d_{1}},\chi_{d_{2}}\}=\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ab}-\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ba}$,
thus $\{\chi_{d_{1}},\chi_{d_{2}}\}\in\operatorname{QInnDer}(\mathbb{C}[G])$. To prove that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$:
$\{\chi_{d_{1}},\chi_{d_{2}}\}(uz,z)=\sum\limits_{ab=zuz^{-1}}\lambda^{a}\mu^{b}-\sum\limits_{ab=u}\lambda^{a}\mu^{b}-\sum\limits_{ba=u}\lambda^{a}\mu^{b}+\sum\limits_{ba=zuz^{-1}}\lambda^{a}\mu^{b}=0$.
We will now turn to the proof of the theorem. Represent $\chi_{d_{0}}$ as
$\chi_{d_{0}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$.
It is sufficient to prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$ using bilinearity of the commutator. Consider the character $\{\chi_{d},\chi^{a}\}$. We need to prove that it is trivial on endomorphisms, i.e. that for all $b\in G$ and for $z\in Z_{G}(b)$ it satisfies $\{\chi_{d},\chi^{a}\}(bz,z)=0$. By Proposition 3.1:
$\{\chi_{d},\chi^{a}\}(bz,z)=\sum\limits_{h\in G}\chi_{d}(bz,h)\chi^{a}(h,z)-\chi^{a}(bz,h)\chi_{d}(h,z)$.
Here $\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and when $h=az$; and $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and when $h=a^{-1}bz$. This means that
$\{\chi_{d},\chi^{a}\}(bz,z)=\chi_{d}(bz,za)-\chi_{d}(bz,az)+\chi_{d}(a^{-1}bz,z)-\chi_{d}(bza^{-1},z)$.
But also
$(a^{-1}bz,z)\circ(bz,za)=(bz,az)\circ(bza^{-1},z)$.
Hence
$\chi_{d}(bz,za)+\chi_{d}(a^{-1}bz,z)=\chi_{d}(bz,az)+\chi_{d}(bza^{-1},z)$,
$\{\chi_{d},\chi^{a}\}(bz,z)=0$.
It is easy to see that central derivations are not quasi-inner by repeating the reasoning from proposition 2.3.
3 Description of the structure of a DG–algebra in terms of characters on a groupoid.
Our next goal is to give a description of the possible DG–algebra structures up to isomorphism. We follow the definitions of [FHT01].
Definition 3.1.
Let $A\cong\bigoplus A_{i}$ be a $\mathbb{Z}$–graded algebra equipped with a derivation
$d:A\rightarrow A$
of degree $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfying the following conditions:
  1. 1.
    $d\circ d=0$,
  2. 2.
    $d(uv)=d(u)v+(-1)^{|u|}ud(v)$.
We say that $(A,d)$ is a differential graded algebra or DG–algebra.
We will give the necessary and sufficient conditions from the point of view of groupoid characters for a derivation to define a DG–algebra structure on a graded group algebra. We will describe the cochain complex case, since the reasoning in the chain complex is similar.
Theorem 3.1.
A character $\chi$ on the groupoid $\Gamma$ defines a DG–algebra structure if and only if the following conditions are satisfied:
  1. 1.
    The convolution is trivial:
    $\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0,\quad\forall g,x\in G.$ (7)
  2. 2.
    For all $(h,g)\in Hom(\Gamma_{A_{i}})$, $i\neq 1$, we have $\chi(h,g)=0$.
The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$ and the second one is equivalent to $d(A^{i})\subseteq A^{i+1}$.
Proof.
Using proposition 2.2, we rewrite the defining properties of a DG–algebra in terms of characters.
  1. 1.
    By direct calculation we can see that
    $d^{2}(g)=d\Big{(}\sum\limits_{h\in G}d^{h}_{g}h\Big{)}=\sum\limits_{h\in G}d^{h}_{g}\Big{(}\sum\limits_{x\in G}d^{x}_{h}x\Big{)}=\sum\limits_{h\in G}(h,g)\Big{(}\sum\limits_{x\in G}(x,h)x\Big{)}=\sum\limits_{x\in G}\Big{(}\sum\limits_{h\in G}(\chi(h,g)\chi(x,h))\Big{)}x.$
  2. 2.
    Let $g\in A^{i}$, then $d(g)=\sum\limits_{h\in G}\chi(h,g)h\in A^{i+1}\Leftrightarrow\chi(h,g)=0$ when $h\notin A^{i+1}$.
    Note that if $u\in A^{i+1},v\in A^{i}$, then for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1}$, $t(\phi)=uv^{-1}\in A^{1}$,which means that in the case of a DG–algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
Now let us recall the definition of an isomorphism of DG–algebras.
Definition 3.2.
An isomorphism $f$: $A_{1}$ $\to$ $A_{2}$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$ is an algebra isomorphism such that following properties hold:
  1. 1.
    f respects the grading, $f(A^{i}_{1})\subset A^{i}_{2}$;
  2. 2.
    for all $a\in A_{1}$, $f(d_{1}(a))=d_{2}(f(a))$.
Now we will describe the structures of DG–algebras up to isomorphism per definition (3.2).
Let $X_{i},i=1,2$ be two matrices (infinity ones, if $G$ is an infinite group) in following sense:
$X_{i}:G\times G\to\mathbb{C},$
$(x,h)\mapsto\chi_{i}(x,h).$
Theorem 3.2.
Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix $C$ such that
  1. 1.
    matrices $X_{1}$ and $X_{2}$ are conjugate, $CX_{2}=X_{1}C;$
  2. 2.
    the mapping f is in $\operatorname{Aut}(\mathbb{C}[G])$.
In other words, that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in\operatorname{Aut}(\mathbb{C}[G])$ and
$\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)=\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h},\forall g,x\in G,$ (8)
where $\chi_{1}$ and $\chi_{2}$ denotes characters that define derivations $d_{1}$ and $d_{2}$.
Proof.
Suppose that we have an isomorphism of DG–algebras $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ ,given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$. By definition, we have
$f(d_{1}(g))=d_{2}(f(g))$.
Then
$f(d_{1}(g))=f\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)h\Big{)}=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)\big{(}\sum\limits_{x\in G}c_{x}^{h}x\Big{)}=\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h}\Big{)}x$;
$d_{2}(f(g))=d_{2}\Big{(}\sum\limits_{h\in G}c_{h}^{g}h\Big{)}=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=\Big{(}\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)\Big{)}x$.
Whence we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, get the necessary equalities. ∎
Corollary 3.1.
In the case when $f\in\operatorname{Aut(G)}$ the equality (8) takes the form
$\chi_{1}(h,g)=\chi_{2}(f(h),f(g)).$
That is, all characters giving isomorphic structures on a group algebra differ by the action of the automorphism $f$ on the groupoid $\Gamma$.
Example 3.1.
In a group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$ such that $A^{1}\neq 0$, and given $a\in Z(G)\cap A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines a DG–algebra structure on $\mathbb{C}[G]$. Note that the identity element of the group lies in $A^{0}$.
Proof.
We have to show that $d^{2}=0$ for such a derivation. We know from Theorem (3.1) that this condition is equivalent to
$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$; $\forall g,x\in G.$
Consider the product $\chi(h,g)\chi(x,h)$. By the definition of an inner derivation, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$; but this is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$; but this is similarly impossible because $|h|\neq|g|.$
The remaining two cases are similar. ∎
Proposition 3.1.
Let $d^{\tau}_{z}(g)=\tau(g)gz$ be the central derivation given by the element $z\in A^{1}$. It defines a DG–algebra structure if and only if $\tau(g)\tau(gz)=0\;\forall g\in G$.
Proof.
The fact that $d^{\tau}_{z}(g):A^{i}\rightarrow A^{i+1}$ follows from the fact that $z\in A^{1}$.
It remains to check under what conditions $d^{\tau}_{z}(g)^{2}=0$. To do this, we simply write by definition:
$(d^{\tau}_{z})^{2}(g)=d^{\tau}_{z}(\tau(g)gz)=\tau(g)\tau(gz)gz^{2}\Rightarrow(d^{\tau}_{z})^{2}(g)=0\Leftrightarrow\tau(g)\tau(gz)=0.$
The example below illustrates a non-trivial condition that is implied by the previous proposition.
Example 3.2.
[…]
defines a DG–algebra structure on $\mathbb{C}[G]$.
[…]
Proof.
It is only necessary to check that the graded Leibniz rule is satisfied. To do this, we can make sure that the given $\tau$ is indeed a group character and then take advantage of Proposition (3.1). It is not difficult to do this by simply going through all the possibilities. ∎
Corollary 3.1.
In the graded group algebra $\mathbb{C}[\mathbb{Z}]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define a DG–algebra structure using the central derivation $d^{\tau}_{x}$, where
[…]
Corollary 3.2.
Consider the Heisenberg integer group of 3×3 upper–triangular matrices
$H_{3}=\left\langle{x,y,z\mid z=xyx^{-1}y^{-1},xz=zx,yz=zy}\right\rangle$.
It is well known that any element of this group can be represented as $y^{b}z^{c}x^{a}$. Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ is an epimorphism into a group of integers. From here we get the grading $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$. And the central derivation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives a DG–algebra structure on $\mathbb{C}[H_{3}]$.